OpenStudy (anonymous):

Given the indicated parts of triangle ABC with y= 90 degrees, approximate the remaining parts. (Round your answers to one decimal place.) a = 26, b = 54 a= b= c=

OpenStudy (anonymous):

b

OpenStudy (kamibug):

It's asking you to solve for the degrees of the other angles or the length of the sides?

OpenStudy (anonymous):

The degrees of a and b but the side of c

OpenStudy (kamibug):

Ok, is there an image?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

Thanks for the help by the way.

OpenStudy (kamibug):

Sure, I'll do my best. :)

OpenStudy (anonymous):

Is it okay if i attach a word document with the Image on it?

OpenStudy (kamibug):

Yeah that's okay

OpenStudy (kamibug):

Okay, so I'm assuming side C is the hypotenuse. To solve for the length of the hypotenuse, you can use the Pythagorean Theorem! :) a^2 + b^2 = c^2 Just plug in the known values for a and b and solve for c. 26^2 + 54^2 = c^2 676 + 2916 = c^2 3592 = c^2 c = \[2\sqrt{898}\]

OpenStudy (kamibug):

To solve for the other angles you can use what you've learned about Since, Cosine, or Tangent. :) Is there a picture of the triangle? That way I'll know the position of each angle and side to be sure.

OpenStudy (anonymous):

how did you get to 2 square root of 898, instead of just a number with a decimal

OpenStudy (anonymous):

no, theres no picture

OpenStudy (kamibug):

Oh, that's because I simplified the square root of 3592. But in decimal form that would be an irrational decimal, but rounds to about 59.93.

OpenStudy (anonymous):

okay, youve been so helpful! I think i will be able to get the other answers.

OpenStudy (anonymous):

is there anyway i can give you good rep? sorry this is my first time using this site. Do i just press best response?

OpenStudy (kamibug):

Lol okay, so I'm just going to guess the triangle looks something like this...